OpenStudy (anonymous):

How do you solve this Exponential equation?: y = 4 x (the "x" is placed as a square root). need help finding the points and other information, so that I can graph. I understand how to graph, im just struggling getting to that step, because I do not know how to even begin this problem.

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

to solve \[y=4^x\] \[x=\frac{\ln(y)}{ln(4)}\]

OpenStudy (anonymous):

oh the x is in a square root! you are supposed to graph \[y=\sqrt{x}\]?

OpenStudy (anonymous):

points on the graph such as (0,0), (1,1), (4,2), (9,3) etc. it is the top half of \[y^2=x\] a parabola with vertex (0,0) facing to the right.

OpenStudy (anonymous):

when you said an exponential equation i thought you meant \[y=4^x\] sorry

OpenStudy (anonymous):

@ satellite73, no you are right that is the right equation

OpenStudy (anonymous):

oh so there is no square root?

OpenStudy (anonymous):

or is it \[y=4^{\sqrt{x}}\]

OpenStudy (anonymous):

\[y = 4^{x}\] is the equation

OpenStudy (anonymous):

i think it is 4 to the x root, y=4^x

OpenStudy (anonymous):

well if is just a matter of graphing \[y=4^{\sqrt{x}}\] pick some nice numbers and plot some points. for example pick x= 0, x = 1, x = 4, x = 9, x = 16 (numbers you know the square root of) and see what you get. (0,0), (1,4), (4,16), (9,64) and so on. you will get a nice exponential graph

OpenStudy (anonymous):

the confusion is the word "solve" which usually means write one variable in terms of another or find its value. i think here you meant "solve the problem"

OpenStudy (anonymous):

here is a nice picture of \[y=4^{\sqrt{x}}\]

OpenStudy (anonymous):

x \ o \ 1 \ 2 \ 3 \ 4 \ 5 \ ----------\------------------------ y=4^(x)\ 1 \ 4 \ 16 \ 64 \256\ 1024 \

OpenStudy (anonymous):

awesome! Thank you satellite73, and mark o. !! you guys are life savers!

OpenStudy (anonymous):

ok wc, good luck